OpenStudy (anonymous):

y=e^-2x+10 as a transformation of y=e^x Write the series of transformation

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

so basically I'm supposed to describe the horizontal shifts and vertical pelletfs keeping in mind of stretches and compressions

OpenStudy (anonymous):

so for instance, y=1/2(x+6)^2-9 would be a horizontal shift left 6 units followed by a vertical compression of 1/2 and then a vertical shift down 9 units

ganeshie8 (ganeshie8):

right, lets start with parent graph y = e^x

ganeshie8 (ganeshie8):

y = e^(2x) compresses the parent graph by 2

ganeshie8 (ganeshie8):

after that, y=e^(-2x) flips the graph vertically

ganeshie8 (ganeshie8):

after that, y = e^(-2x) + 10 shifts the graph up by 10 units

OpenStudy (anonymous):

okay so that means there is a vertical compression of 2 and a reflection across the x axis and a vertical shift up 10 unites

ganeshie8 (ganeshie8):

y = e^(2x) compresses the parent graph `horizontally` by 2

OpenStudy (anonymous):

oh right

ganeshie8 (ganeshie8):

okay so that means there is a `horizontal` compression of 2 and a reflection across the `y` axis and a vertical shift up 10 unites

ganeshie8 (ganeshie8):

see if that looks good ^

OpenStudy (anonymous):

yep. i meant to type y